Direction:Given is the selected information of Alpha Corporation. Answer the questions that 

follow. Show your computation.

December 31, 2019

December 31, 2020

Cash 

10

,

000

80

,

000

Marketable Securities   

30

,

000

10

,

000

Accounts Receivable (Net)

50

,

000

150

,

000

Inventories

90

,

000

150

,

000

Land and Building (Net)

340

,

000

360

,

000

Notes Payable, Trade

70

,

000

110

,

000

Accounts Payable, Trade

20

,

000

40

,

000

Mortgage Payable

270

,

000

280

,

000

Cash Sales

1,800,000

1,600,000

Credit Sales 

500,000

800,000

Cost of Sales

1,000,000

1,400,000

1. Alpha’s Acid test ratio as of December 31, 2020.

2. Alpha’s current ratio as of December 31, 2020.

1. Alpha’s receivable turnover as of December 31, 2020.

1. Alpha’s inventory turnover as of December 31, 2020.


1
Expert's answer
2021-06-07T14:12:17-0400

1.

Acid test ratio"=\frac{Current\ Assets-Inventory}{Current\ Liabilities}"


Assets: Cash, Land and Building, Accounts Receivable, Cash Sales, Credit Sales

Liabilities: Accounts Payable, Notes Payable, Mortgage Payable, Cost of Sales


For December 31, 2020:

Assets"=80000+360000+150000+1600000+800000=3140000"

Liabilities="40000+280000+110000+1400000=1830000"

Inventories"=150000"


Acid test ratio"=\frac{3140000-150000}{1830000}=1.63"


2.

current ratio="\frac{Current\ Assets}{Current\ Liabilities}=\frac{3140000}{1830000}=1.72"


3.

receivable turnover"=(Net\ credit\ sales)/(net\ accounts\ receivable)"


"Net\ credit\ sales=\frac{500000+800000}{2}=650000"

"net\ accounts\ receivable=\frac{55000+150000}{2}=200000"


receivable turnover="650000\/200000=3.25"


4.

inventory turnover"=\frac{nrt\ sales}{average\ inventory}=\frac{1200000}{120000}=10"